How many numbers each less than 400 can be formed from the digits 1,2,3,4,5,6, if repetition is allowed?
Answer by checkley71(8403)

3 DIGIT NUMBERS=3*6*6=108
2 DIGIT NUMBERS=6*6=36
1 DIGIT NUMBERS=6
108+36+6=150 TOTAL NUMBERS OF NUMBERS < 400.
EXAMPLES
111 TO 366
11 TO 66
1 TO 6